### v1

@@ 5.1 Task
We selected seven images from Where’s Wally. An image, featuring the fewest characters and minimal distractions, was used for training so that the participants could familiarize themselves with the three interaction scenarios (Static, Parallel, and Tilt). The remaining six images were classified into three low- and three medium-difficulty categories according to difficulty based on approximate character count and background complexity.
@@ 5.2 Task Image Characteristics and Visual Challenge
The pictures selected for the experiment came from the ”Where’s Wally?”222Where’s Wally? Images © Martin Handford / Walker Books Ltd. Used under fair use for academic research purposes. All images used in the study, including training and trial stimuli, are available in the supplementary materials accompanying this article. series and varied considerably in visual complexity.
@@
The training image, city, is the least complex of the set, featuring a modest number of characters with relatively clear segmentation and minimal foreground clutter. In contrast, the remaining six images vary widely in layout and visual load. Others, e.g. beach, introduce dense and unstructured clustering of figures, many partially occluded by umbrellas and beach props, which creates substantial camouflage effects. Other images are highly populated but exhibit more regular spatial organization.
@@
5.3 Prototype Design
We developed our task prototype in Unity3D (version 6) for the Meta Quest 3 VR HMD. As shown in HeadZoom: Hands-Free Zooming and Panning for 2D Image Navigation Using Head Motion.c, each trial presented participants with a large Where’s Wally image (2800×1749 pixels). A red ring (approximately 105 pixels in radius) overlaid the image to indicate the current raycast hit point. In the upper left corner of the image, the participants saw the assigned interaction scenario (e.g. static, parallel, or tilt) and the trial number, ensuring that they were aware of the condition they were experiencing.
@@ 5.4 Experimental Design
Our study adopted a within subjects design for the participants that included three interaction scenarios (Static, parallel, and tilt) and two difficulty levels (low and medium). Figure 4.a shows an overview of the experimental timeline. Each participant was randomly assigned one low and one medium difficulty image under each scenario, resulting in six total trials (3 scenarios x 2 difficulty levels). We counterbalanced and randomized the order of both scenarios and difficulty levels to mitigate the potential bias from time‐sequence effects or subjective judgments.
Each scenario consisted of two trials (one low-difficulty image and one medium-difficulty image), presented in a counterbalanced sequence. Upon finishing each trial, participants completed a subjective rating (on a 1–7 scale) to indicate their perceived challenge in locating Wally within that image.
After completing both trials in a scenario, participants removed the headset and filled a post-scenario questionnaire assessing system usability and user experience. Upon concluding all three scenarios (i.e., six trials), they completed a final post-experience questionnaire, ranking each interaction technique in terms of zooming efficiency and object-finding accuracy.
@@ 5.5 Procedure
Figure 4.b illustrates the overall experimental setup. Upon arrival, participants were given an information sheet describing the procedures, followed by completion of a consent form and a pre‐experience questionnaire. Then, we configured the Quest headset and seated participants in a fixed‐position chair (Figure 4.c). We used Meta Horizon to stream the headset’s view for monitoring participant performance. Participants then underwent a 15 minute training session, familiarising themselves with all three interaction scenarios for the Where’s Wally task. Participants then proceeded according to the process depicted in 5.4. We used Google Forms to collect all pre-, post‐scenario, and post‐experience responses. The experiment lasted approximately 40 minutes, after which each participant received a $20 gift card as compensation.
@@ 5.6 Participants
We recruited 31 participants (16 women, 15 men, ages 19 to 49). A total of 8 participants (10. 5%) reported never using VR, 19 participants (61.3%) had fewer than 5 hours of VR experience, 1 (3.2%) reported having 5–20 hours of experience, 8 participants (25.8%) had logged more than 20 hours, and 3 participants (9.7%) had never used VR. Among all participants, 8 (25.8%) indicated they had never experienced VR‐related motion sickness, 4 (12.9%) were unsure or had insufficient VR experience, and 19 (61.3%) experienced it only occasionally.
@@ 6.2 Spatial Metrics
Total Head Movement. Is is the cumulative headset displacement where represents headset position at frame .:
Total Head Rotation. Summation of angular displacement between consecutive forward vectors :
@@
Hover Time on Characters. Time spent, gazing near a specific character (Wally, Wenda, Wizard, Odlaw):
with binary indicator:
@@
Average Zoom indicates participants’ average zoom level throughout the trial. Maximum Zoom records the highest level of magnification achieved during the trial.
@@ 6.4 Subjective Difficulty Metrics
Participants rated each image’s difficulty using a 1–7 scale (ImageUserGrade), where 1 = ”very difficult” and 7 = ”very easy.” This subjective assessment allows correlation analysis with objective performance metrics (e.g., Completion Time, False Positives), highlighting discrepancies or consistencies between perceived difficulty and actual performance.
@@ 7 Results
To evaluate the performance and interaction characteristics of the three interaction techniques (Static, Parallel, and Tilt), we performed a comprehensive statistical analysis on data collected from a controlled VR experiment involving 31 participants. The task involved participants in locating a target character (Wally) within images of varying complexity. The statistical evaluation process consisted of three main stages and is presented in Figure 5:
@@ 7.1 Interaction Pattern Differences
The ANOVA results revealed significant differences among interaction modes for out of metrics analyzed, with F-statistics ranging from to (all ). The most pronounced differences were observed for Average Head Rotation (, ) and Average Head Movement (, ).
After the Bonferroni correction, out of pairwise comparisons remained statistically significant. The key differences were:
@@
Tilt Mode generally showed intermediate values for most physical metrics, indicating moderate physical interaction compared to the other two modes.
Figure 7 shows the normalized averaged performance metrics for Total Head Movement, Max Zoom, Total Zoom Distance, Zoom Change Count and Max Lean, revealing that Parallel is the most ergonomic mode, since it leads to lower physical movements. Figure 8 presents the distributions of the Total Head Rotation per interaction mode, indicating that the Static mode required more movements when compared to the other interaction modes.
@@ 7.2 Task Performance and Efficiency
Despite significant differences in physical interaction patterns, no statistically significant differences emerged regarding task completion time (, ). Mean completion times were similar across modes: Static with secs (), Parallel with seconds (), and Tilt with seconds (). Additionally, no significant differences were found for false positive rates or subjective difficulty ratings. This suggests that all three interaction techniques provided similar efficiency and accuracy, despite differences in physical interaction demands. Figure 6 shows the distribution of the completion times across participants.
@@ 7.3 Usability Evaluation
The System Usability Scale (SUS) scores, assessing subjective usability, did not show significant differences between interaction modes (). All three modes achieved high usability scores: Static (), Parallel (), and Tilt ().
These scores indicate excellent usability for all interaction methods. The analysis of individual SUS items also revealed that there were no significant differences in perceived complexity, ease of use, learnability, confidence, or support needs, confirming that differences in physical interaction did not significantly influence overall usability perceptions.
@@ 7.4 User Preferences
The Parallel mode was most frequently selected as the preferred method, with the highest ”best” ratings, while Static and Tilt modes were more frequently identified as the least preferred. Specifically, Parallel was chosen most frequently as the ”best” interaction mode by participants and less often marked as the ”worst,” contrasting with Tilt, which received the highest number of ”worst” ratings. This subjective preference aligns with objective findings on reduced physical demand for the Parallel mode, underscoring user appreciation for its ergonomic efficiency.
@@ 7.5 Impact of Image Complexity
“Time On” metrics (TimeOnWally, TimeOnWenda, TimeOnWizard, TimeOnOdlaw) indicate the duration participants focused on specific targets or distractors did not differ significantly across interaction modes. This finding held across varying image complexities, suggesting consistent interaction patterns regardless of visual clutter.
@@ 7.6 Gender Differences
In terms of gender, only the User Difficulty Rating showed significant gender differences, with males rating tasks as slightly easier () compared to women (). Further analysis revealed a significant gender-interaction mode effect for SUS scores specifically in the Parallel mode, where males reported significantly higher usability scores () compared to females (), indicating gender may influence subjective experience particularly in this interaction condition.
@@ 8 Discussion
Finding 1: Significant Differences in Physical Interaction Patterns. Our study revealed significant differences in physical interaction patterns among the three modes (Static, Parallel, and Tilt), particularly regarding head movement, rotation, and zoom usage. The Static mode had significantly higher physical engagement, evidenced by greater head movement and rotation, than the Parallel mode. Conversely, the Parallel mode demonstrated minimal physical exertion requirements, making it the most ergonomically efficient. This aligns with the participants’ subjective preferences, as the Parallel mode was frequently the preferred interaction mode and less frequently the least preferred. The Tilt mode generally showed moderate values, indicating that this technique balances physical engagement and comfort. This finding is important, as it suggests that interaction techniques significantly affect user comfort and could influence user adoption, especially in prolonged VR applications or precision-demanding tasks like medical imaging analysis.
@@
Finding 2: Comparable Task Performance Across Interaction Modes. Despite pronounced physical differences among interaction techniques, no significant differences were observed in task completion times, accuracy, or subjective difficulty ratings. This indicates that while the techniques differ significantly in ergonomic efficiency, they do not affect the effectiveness or accuracy of task execution. These results suggest that designers of VR interfaces have flexibility in choosing interaction techniques based on considerations other than raw performance metrics, such as user comfort, accessibility, or specific application requirements. Thus, the decision regarding interaction technique selection can prioritize user comfort and reduce physical strain without compromising task performance.
@@
Finding 3: High Usability Scores Independent of the Interaction Technique. All interaction modes received high usability ratings, with no significant differences observed in the overall SUS scores or in individual usability dimensions such as complexity, ease of use, learnability and confidence. This suggests that despite differences in how users physically interacted with the VR system, all interaction methods provided an excellent user experience. High usability scores across all modes imply that users readily adapted to different interaction paradigms, reinforcing that usability perceptions are robust to variations in physical interaction patterns. Consequently, one can confidently employ varied interaction methods tailored to specific contextual or ergonomic considerations without concern for negative usability impacts.
@@
Finding 4: Gender Differences in Interaction Experiences. Our analysis identified notable gender-related differences, specifically in subjective task difficulty ratings and usability perceptions within the Parallel interaction mode. Male participants rated the tasks as slightly easier and reported significantly higher usability for the Parallel mode than female participants. These gender differences highlight the importance of considering demographic factors in interaction design, suggesting that user groups might perceive certain interaction methods differently.
@@
Finding 5: Robustness of Observed Differences. Applying the Bonferroni correction minimally impacted the significant differences observed in our analysis, with only a minor reduction from 16 to 15 significant pairwise comparisons. This highlights the robustness of our findings, indicating that the reported differences in physical interaction patterns and ergonomic efficiency are reliable and unlikely to be Type I errors. The large effect sizes observed, particularly between Static and Parallel modes, further substantiate the practical significance of these differences. These robust results reinforce the recommendation of the Parallel mode as an ergonomic choice for extended or precision-based VR tasks.
@@
Our findings underscore the importance of ergonomic considerations in VR interaction design, demonstrate flexibility in achieving comparable performance outcomes across interaction methods, and highlight user preferences and demographic differences as critical considerations for user-centered VR system development.
@@ 9 Design Guidelines
Based on the user study results and our findings, there are some guidelines that could be provided for people developing hands-free, controller-free VR interaction, including:
G1: Minimize effort: User interface techniques should be developed that favor mappings with low head movement requirements (e.g., Parallel) to reduce fatigue.
G2: Stabilize input with filtering: Use smoothing filters to suppress jitter in continuous control tasks.
G3: Ensure spatial consistency: Maintain predictable image behavior to reduce disorientation (as seen in Tilt).
G4: Support Visual Anchoring: Design interactions that allow users to maintain visual focus during input. Avoid mappings that disrupt line-of-sight or require large shifts in viewpoint.
G5: Preserve visual focus: Avoid techniques that disrupt gaze alignment during navigation.
G6: Balance control and simplicity: Choose interaction modes that generalize well and require minimal learning.

### v2

@@ 4.1 Task
We selected seven images from Where’s Wally. An image, featuring the fewest characters and minimal distractions, was used for training so that the participants could familiarize themselves with the three interaction modes (Static, Parallel, and Tilt). The remaining six images were classified into two difficulty levels (three easy and three hard) according to difficulty based on approximate character count and background complexity.
@@ 4.2 Task Image Characteristics and Visual Challenge
The pictures selected for the experiment came from the ”Where’s Wally?”222Where’s Wally? Images © Martin Handford / Walker Books Ltd. Used under fair use for academic research purposes. All images used in the study, including training and trial stimuli, are available as supplementary materials to this article. series and varied considerably in visual complexity.
@@
The training image, city, is the least complex of the set, featuring a modest number of characters with relatively clear segmentation and minimal foreground clutter. In contrast, the remaining six images vary widely in layout and visual load. Others, e.g. beach, introduce dense and unstructured clustering of figures, many partially occluded by umbrellas and beach props, creating substantial camouflage effects. Other images are highly populated but exhibit more regular spatial organization.
@@
4.3 Task Prototype Design
We developed our task prototype in Unity3D (version 6) for the Meta Quest 3 VR HMD. As shown in HeadZoom: Hands-Free Zooming and Panning for 2D Image Navigation Using Head Motion.c, each trial presented participants with a large Where’s Wally image (2800×1749 pixels). A red ring (approximately 105 pixels in radius) overlaid the image to indicate the current raycast hit point. In the upper left corner of the image, the participants saw the assigned interaction mode (e.g. static, parallel, or tilt) and the trial number, ensuring that they were aware of the condition they were experiencing.
@@ 4.4 Experimental Design
Our study adopted a within subjects design for the participants that included three interaction modes (Static, parallel, and tilt) and two difficulty levels (easy and hard). Figure 3.a shows an overview of the experimental timeline. Each participant was randomly assigned one low and one hard difficulty image under each mode, resulting in six total trials (3 modes x 2 difficulty levels). We counterbalanced and randomized the order of both mode and difficulty levels to mitigate the potential bias from time‐sequence effects or subjective judgments.
Each condition consisted of two trials (one low-difficulty image and one high-difficulty image), presented in a counterbalanced sequence. Upon finishing each trial, participants completed a subjective rating (on a 1–7 scale) to indicate their perceived challenge in locating Wally within that image.
After completing both trials in a scenario, participants removed the headset and filled a post-condition questionnaire assessing system usability and user experience. Upon concluding all three modes (i.e., six trials), they completed a final post-experience questionnaire, ranking each interaction technique in terms of zooming efficiency and object-finding accuracy.
@@ 4.5 Procedure
Figure 3.b illustrates the overall experimental setup. Upon arrival, participants were given an information sheet describing the procedures, followed by completion of a consent form and a pre‐experience questionnaire. Then, we configured the Quest headset and seated participants in a fixed‐position chair (Figure 3.c). We used Meta Horizon to stream the headset’s view for monitoring participant performance. Participants then underwent a 15 minute training session, familiarising themselves with all three interaction modes for the Where’s Wally task. Participants then proceeded according to the process depicted in 4.4. We used Google Forms to collect all pre-, post‐condition, and post‐experience responses. The experiment lasted approximately 40 minutes, after which each participant received a $20 gift card as compensation.
@@ 4.6 Participants
We recruited 31 participants (16 women, 15 men, ages 19 to 49). A total of 8 participants (25.8%) reported never using VR, 19 participants (61.3%) had fewer than 5 hours of VR experience, 1 (3.2%) reported having 5–20 hours of experience, 8 participants (25.8%) had logged more than 20 hours, and 3 participants (9.7%) had never used VR. Among all participants, 8 (25.8%) had never experienced VR‐related motion sickness, 4 (12.9%) were unsure or had insufficient VR experience, and 19 (61.3%) experienced it only occasionally.
@@ 5.2 Spatial Metrics
Total Head Movement. It is the cumulative headset displacement where represents headset position at frame .: .
Total Head Rotation. Summatin of angular displacement between consecutive forward vectors : .
@@
Hover Time on Characters. Time spent, gazing near a specific character (Wally, Wenda, Wizard, Odlaw) with binary indicator .
@@
Average Zoom indicates participants’ average zoom level throughout the trial. Similar definitions apply for AvgHeadRotation and AvgHeadMovement. Maximum Zoom records the highest level of magnification achieved during the trial.
@@ 5.4 Subjective Difficulty Metrics
Participants rated each image’s difficulty using a 1–7 scale (ImageUserGrade), where 1 = “very difficult” and 7 = “very easy”. This allows correlation analysis with objective performance metrics (e.g., Completion Time, False Positives) showing discrepancies or consistencies between perceived difficulty and actual performance.
@@ 6 Results
To evaluate the performance and interaction characteristics of the three interaction techniques (Static, Parallel, and Tilt), we performed a comprehensive statistical analysis on data collected from a controlled VR experiment involving 31 participants. The task involved participants in locating a target character (Wally) within images of varying complexity. The statistical evaluation process consisted of three main stages and is presented in Figure 4.
@@ 6.1 Interaction Pattern Differences
The ANOVA results revealed significant differences among interaction modes for out of metrics analyzed, with F-statistics ranging from to (all ). Effect sizes ranged from to . The most pronounced differences were observed for Average Head Rotation (, , ) and Average Head Movement (, , ). After the Bonferroni correction (), out of pairwise comparisons remained statistically significant. Post-hoc -statistics ranged from to , with degrees of freedom varying from to depending on equality of variances, and Cohen’s effect sizes ranging from to . The key differences were:
@@
Tilt Mode generally showed intermediate values for most physical metrics, indicating moderate physical interaction compared to the other two modes. Figure 7 shows the normalized averaged performance metrics for Total Head Movement, Max Zoom, Total Zoom Distance, Zoom Change Count and Max Lean, revealing that Parallel is the physically less demanding mode, as it leads to lower physical movements. Figure 6 presents the distributions of the Total Head Rotation per interaction mode, indicating Static mode required more movements when compared to the other interaction modes.
@@ 6.2 Task Performance and Efficiency
Despite significant differences in physical interaction patterns, no statistically significant differences emerged regarding task completion time (, ). Mean completion times were similar across modes: Static with secs (), Parallel with seconds (), and Tilt with seconds (). Additionally, no significant differences were found for false positive rates (, ) or subjective difficulty ratings (, ). This suggests that all three interaction techniques provided similar efficiency and accuracy, despite differences in physical interaction demands. Figure 5 shows the distribution of the completion times across participants.
@@ 6.3 Usability Evaluation
The System Usability Scale scores, assessing subjective usability, did not show significant differences between interaction modes (). All three modes achieved high usability scores: Static (), Parallel (), and Tilt (). These scores indicate excellent usability for all interaction methods. The analysis of individual SUS items also revealed that there were no significant differences in perceived complexity, ease of use, learnability, confidence, or support needs, confirming that differences in physical interaction did not significantly influence overall usability perceptions.
@@ 6.4 User Preferences
The Parallel mode was most frequently selected as the preferred method, with the highest “best” ratings (13 participants, 41.94%), while Static and Tilt modes were more frequently identified as the least preferred. Specifically, Parallel was chosen most frequently as the “best” interaction mode by participants and less often marked as the “worst” (6 participants, 19.35%), contrasting with Tilt, which received the highest number of “worst” ratings (13 participants, 41.94%). Static mode received intermediate preference ratings with 8 participants (25.81%) rating it as best and 12 participants (38.71%) rating it as least preferred. This subjective preference aligns with objective findings on reduced physical demand for the Parallel mode, underscoring user appreciation for its efficiency.
@@ 6.5 Impact of Image Complexity
“Time On” metrics (TimeOnWally, TimeOnWenda, TimeOnWizard, TimeOnOdlaw) indicate the duration participants focused on specific targets or distractors did not differ significantly across interaction modes. Specifically, TimeOnWaldo (), TimeOnWenda (), TimeOnWizard (), and TimeOnOdlaw () all showed non-significant effects. However, Figure 8 reveals substantial variability in success rates across images and interaction modes, with some images showing large differences (e.g., Medieval: to , Ski: to ). This reflects inherent differences in difficulty in Where’s Waldo images, where visual complexity and target detectability vary systematically across scenes. This finding held across varying image complexities, suggesting consistent interaction patterns regardless of visual clutter.
@@ 6.6 Gender Differences
In terms of gender, only the User Difficulty Rating showed significant gender differences, with males rating tasks as slightly easier () compared to women (; ). Further analysis revealed a significant gender-interaction mode effect for SUS scores specifically in the Parallel mode, where males reported significantly higher usability scores () compared to females (; ), indicating gender may influence subjective experience particularly in this interaction condition.
@@ 7 Discussion
Finding 1: Significant Differences in Physical Interaction Patterns. Our study revealed significant differences in physical interaction patterns among the three modes (Static, Parallel, and Tilt), particularly regarding head movement, rotation, and zoom usage. The Static mode had significantly higher physical engagement, evidenced by greater head movement and rotation, than the Parallel mode. Conversely, the Parallel mode demonstrated minimal physical exertion requirements, making it the most physically less demanding. This aligns with the participants’ subjective preferences, as the Parallel mode was frequently the preferred interaction mode and less frequently the least preferred. The Tilt mode generally showed moderate values, indicating that this technique balances physical engagement and comfort. This finding is important, as it suggests that interaction techniques significantly affect user comfort and could influence user adoption, especially in prolonged VR applications or precision-demanding tasks like medical imaging analysis.
@@
Finding 2: Comparable Task Performance Across Interaction Modes. Despite pronounced physical differences among interaction techniques, no significant differences were observed in task completion times, accuracy, or subjective difficulty ratings. This indicates that while the techniques differ significantly in physical demand, they do not affect the effectiveness or accuracy of task execution. These results suggest that designers of VR interfaces have flexibility in choosing interaction techniques based on considerations other than raw performance metrics, such as user comfort, accessibility, or specific application requirements. Thus, the decision regarding interaction technique selection can prioritize user comfort and reduce physical strain without compromising task performance.
@@
Finding 3: High Usability Scores Independent of the Interaction Technique. All interaction modes received high usability ratings, with no significant differences observed in the overall SUS scores or in individual usability dimensions such as complexity, ease of use, learnability and confidence. This suggests that despite differences in how users physically interacted in VR, all methods provided an excellent user experience. High usability scores across all modes imply that users readily adapted to different interaction paradigms, reinforcing that usability perceptions are robust to variations in physical interaction patterns. Consequently, one can confidently employ varied interaction methods tailored to specific contextual considerations without concern for negative usability impacts.
@@
Finding 4: Gender Differences in Interaction Experiences. Our analysis identified gender-related differences, specifically in subjective task difficulty ratings and usability perceptions within the Parallel interaction mode. Males rated the tasks as slightly easier and reported significantly higher usability for the Parallel mode than females. These gender differences highlight the importance of considering demographic factors in interaction design, suggesting that user groups might perceive certain interaction methods differently.
@@
Finding 5: Robustness of Observed Differences. Applying the Bonferroni correction minimally impacted the significant differences observed, with only a minor reduction from 16 to 15 significant pairwise comparisons. This highlights the robustness of our findings, indicating that the reported differences in physical interaction patterns and physical demand are reliable and unlikely to be Type I errors. The large effect sizes observed, particularly between Static and Parallel modes, further substantiate the practical significance of these differences. These results reinforce the recommendation of the Parallel mode choice for extended or precision-based VR tasks.
@@
Our findings underscore the importance of physical demand considerations in VR interaction design, demonstrate flexibility in achieving comparable performance outcomes across interaction methods, and highlight user preferences and demographic differences as critical considerations for user-centered VR system development.
8 Application Scenarios
Although the primary evaluation of HeadZoom focused on a controlled image search task, the system was designed with broader use cases in mind. To demonstrate the system’s generalizability, we explored how HeadZoom could be applied across several domains requiring precise, hands-free navigation of two-dimensional content.
Geographic Information System (GIS): GIS interfaces rely on frequent panning and zooming using mouse-based interaction (desktop), or multitouch gestures (mobile) [74, 17]. With HeadZoom, users can manipulate tiled map data by leaning forward and backward to zoom in or out, and by rotating their head to pan horizontally or vertically. Users are tasked with locating iconic landmarks in globally recognized cities using only head-based input: leaning forward and backward adjusts the zoom level, while head rotations pan the map horizontally and vertically (see Figure 9.a). We adopted a ‘focus zoom’ approach inspired by focus-plus-context methods [4]: during zoom-out, the area around the user’s gaze retains detail, while the periphery zooms out, reducing abrupt changes and helping users maintain context during navigation.
Medical Imaging: Radiologists navigate large and detailed images, but traditional input devices such as mice can be inefficient. Alternatives such as head tracking, gesture control, and voice commands have been evaluated for image navigation [10, 69]. HeadZoom can be adapted for radiological image inspection, using high-resolution chest X-rays obtained, for example, as 12-bit DICOM files from the MIMIC data set; cf. Figure 9.b. This use case targets professional radiologists and shifts the task paradigm from target localization to clinical evaluation, similarly to [62]. Navigation across image sets is achieved using a wireless button, while a Bluetooth mouse allows users to adjust windowing parameters critical for diagnostic visibility, although these tasks could conceivably be performed via speech actions.
Digital Photography: Photo browsing is typically based on scrolling grids and pinch-to-zoom gestures. Some researchers investigated alternative approaches such as PhotoMesa, a zoomable image browser [5]. HeadZoom enables users to explore large photo collections using head-based zoom and pan controls. Unlike GIS and medical imaging setups, which prioritize analytical precision, this scenario emphasizes natural navigation and user experience. The interface presents image grids where users can seamlessly zoom in individual photos and pan across the collection cf. Figure 9.c, offering a hands-free alternative to traditional gesture-based browsing.
@@ 9 Design Guidelines
Based on our findings, some guidelines could be provided for people developing hands-free, controller-free VR interaction, including:
G1: Minimize effort: (From reduced head movement in Parallel mode). Parallel mode significantly reduced total head movement, minimizing physical effort/fatigue.
G2: Stabilize input with filtering: (Based on our Kalman filter implementation). Kalman filters smoothed head tracking, reducing jitter and improving control precision.
G3: Ensure spatial consistency: (Justified by disorientation in Tilt mode). Tilt mode’s dynamic reorientation was disorienting, highlighting the need for predictable image behavior.
G4: Support Visual Anchoring: (Following slightly faster Static mode). While not significant, Static mode’s faster completion times suggest a stable visual anchor aids visual search. Design interactions that allow users to maintain visual focus during input. Avoid mappings that disrupt line-of-sight or require large shifts in viewpoint.
G5: Balance control and simplicity: (Justified by high SUS scores). All modes achieved high SUS scores ( 80), indicating ease of use.